OpenStudy (vickyisthesmartone123):

The volume of a cube is 400 m3. Which statements are true about the same cube measured in cubic centimeters? Choose all answers that are correct. A. The number representing the volume would decrease. B. The number representing the volume would increase. C. The actual volume would decrease. D. The actual volume would increase. E. The actual volume would stay the same.

OpenStudy (vickyisthesmartone123):

@jacobrocket173

OpenStudy (dmndlife24):

B. and E. i believe

OpenStudy (anonymous):

1 \ m^3=1000000 \ cm^3 \\ \\
400 \ m^3=400 \times 100 00 00 \ cm^3=400000000 \ cm^3 As you can see, the number representing the volume increased. We converted one unit to another but the actual value is the same - 400000000 cm³ is the same as 400 m³. The correct answers are B and E.

OpenStudy (vickyisthesmartone123):

are yu suryodnt get tht fo te intrnet

OpenStudy (anonymous):

no I was solving it in paper

OpenStudy (vickyisthesmartone123):

ok whatever can you help with 4 or eplz?

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

4 more only right